OpenStudy (anonymous):

Which ordered pairs from the replacement set are solutions to the equation? y = -2x + 9 ; Replacement Set {((-1, 11), (0, 8), (1, 7), (2, 5)} can you help me work my way to the answer

OpenStudy (anonymous):

plug in x and y and if they make the equation true then theyre solutions first one: 11=-2(-1)+9 11=2+9 11=11 thats true so first one is a solution, do it with the rest

OpenStudy (anonymous):

thank you